# Scalability of Metropolis-within-Gibbs schemes for high-dimensional Bayesian models

### v1

@@ Abstract
We study general coordinate-wise MCMC schemes (such as Metropolis-within-Gibbs samplers), which are commonly used to fit Bayesian non-conjugate hierarchical models. We relate their convergence properties to the ones of the corresponding (potentially not implementable) Gibbs sampler through the notion of conditional conductance. This allows us to study the performances of popular Metropolis-within-Gibbs schemes for non-conjugate hierarchical models, in high-dimensional regimes where both number of datapoints and parameters increase. Given random data-generating assumptions, we establish dimension-free convergence results, which are in close accordance with numerical evidence. Applications to Bayesian models for binary regression with unknown hyperparameters and discretely observed diffusions are also discussed. Motivated by such statistical applications, auxiliary results of independent interest on approximate conductances and perturbation of Markov operators are provided.
@@ 1 Introduction
Over 30 years ago, the Gibbs sampler (GS) and more general coordinate-wise samplers (often termed Metropolis-within-Gibbs (MwG) samplers) were introduced as powerful tools to enable Bayesian inference for structured data (including spatial, hierarchical or temporal models), for example see [21, 63, 5, 12, 10]. The substantial impact of these methods has spread across a wide variety of scientific fields, for example see [22, 26, 41].
Coordinate-wise Markov chain Monte Carlo (MCMC) schemes work by partitioning the vector of parameters in different blocks and updating them one at the time, conditional on all the others. Working in such a coordinate-wise manner can be computationally beneficial in many cases (see e.g. Section 1.1) and it has been observed empirically that such samplers often have extremely good convergence properties. However theoretical understanding of this phenomenon has proved difficult, as the analysis of the resulting algorithms is often subtle and case-specific. Substantial progress has now been made on understanding the pure GS (see Section 1.3), but implementation of these algorithms is generally restricted to contexts with specialised conditional conjugacy properties and therefore it is important to understand more general coordinate-wise samplers which remain comparatively under-studied.
@@ 1.1 Motivating example: non-conjugate hierarchical models
with $y\in\{0,\dots,m\}$ and $m$ being a positive integer. Under model (1), the conditional distribution of $\bm{\theta}$ given $(Y_{1:J},\mu,\tau)$ factorizes as $\mathcal{L}(\text{d}\bm{\theta}|Y_{1:J},\mu,\tau)=\otimes_{j=1}^{J}\mathcal{L}(\text{d}\theta_{j}|Y_{j},\mu,\tau)$, where $\otimes$ denotes the product of independent distributions. This makes model (1) particularly well-suited for coordinate-wise samplers, since the high-dimensional update of $\bm{\theta}$ given $(\mu,\tau)$ decouples into $J$ independent low-dimensional updates (see Section 6.4 for more discussion on the computational implications of this). Figure 1 compares the efficiency of the resulting samplers with two gradient-based MCMC methods. The target is the joint posterior distribution $\mathcal{L}(\text{d}\bm{\theta},\text{d}\mu,\text{d}\tau|Y_{1:J})$, and we consider the high-dimensional regime where $J\to\infty$, so that both the number of datapoints and parameters, i.e. $n=Jm$ and $p=J\ell+D$ respectively, diverge. Full details on the simulation set-up, including algorithmic and prior specifications, are postponed to Section 6.5.
@@ 1.2 Objective and structure of the paper
The structure of the paper is as follows. Section 2 briefly recalls the notion of conductance and its connection to mixing times of Markov chains. Section 3 introduces the class of coordinate-wise schemes that we consider (which include random-scan GS and MwG) and provides general results relating the conductance of coordinate-wise schemes to the one of GS (Theorem 1 and Corollary 1). Section 4 consider the specific case of MwG schemes and the example of conditionally log-concave targets. After discussing independent results about approximate conductances and perturbations of Markov operators (Section 5), we move to statistical applications, where we consider Bayesian hierarchical models (Section 6), logistic regression with unknown hyperparameters (Section 7) and inference for discretely observed diffusion processes (Section 8). Mathematical proofs, together with some additional examples and results, are postponed to the Appendix.
@@ 1.3 Related literature
Compared to other MCMC schemes (such as gradient-based ones), there are relatively few quantitative theoretical results for coordinate-wise sampling schemes. Moreover, most works focus on the exact Gibbs sampler, e.g. [62, 1, 59, 36, 32, 70, 51, 29, 52, 55, 13], for which specific techniques exploiting the exact updating structure and the associated alternating projection representation can be exploited [58, 15, 16, 3]. Results for general coordinate-wise MCMC, including MwG, are instead quite limited. Exceptions include [56, 42, 30, 31, 53], which mostly focus on geometric or uniform ergodicity, and [67], which provides some analysis of MALA-within-Gibbs schemes.
While working on our manuscript, it came to our attention that [54] independently and concurrently developed results relating the spectral Gap of coordinate-wise samplers to the ones of exact GS. Their results are similar in spirit to the ones we develop in Sections 3 and 4. On the other hand, the specific results and the type of applications are significantly different. For example, [54] study the spectral gap, while in this work we work with the approximate version of the conductance defined in (4), which is crucial for the subsequent applications we consider (see e.g. Remarks 8 and 11). Also, we consider and develop in some detail various statistical applications (Sections 6, 7 and 8) where we combine our techniques with posterior asymptotics, dimensionality reduction and perturbation arguments.
@@ 2 Mixing times and conductance
In order to give quantitative bounds on (3) a typical strategy is to study the $s$-conductance of $P$, i.e.
@@
Let $P$ be a $\pi$-invariant Markov transition kernel. For every $s\in[0,1/2)$, $t\geq 0$, $M\geq 1$ and $\mu\in\mathcal{N}\left(\pi,M\right)$, it holds
@@
The usual definition of $s$-conductance (see e.g. [39, 18]) is slightly different with respect to the one given in (4). Appendix B provides results for the alternative definition. Definition (4) is less tight but leads to neater formulas (see e.g. Theorem 1).
@@ 3 Coordinate-wise MCMC
where $P_{i}^{\textbf{x}_{-i}}$ is a $\pi_{i}(\cdot|\textbf{x}_{-i})$-invariant Markov kernel on $\mathcal{X}_{i}$ and thus, by construction, $P_{i}$ is invariant with respect to $\pi$. At each iteration, a Markov chain evolving according to $P$ picks uniformly at random a coordinate $i$ in $\{1,\dots,d\}$ and then updates $x_{i}$ sampling from $P_{i}^{\textbf{x}_{-i}}(x_{i},\cdot)$, while leaving $\textbf{x}_{-i}$ unchanged. We will sometimes assume each $P_{i}^{\textbf{x}_{-i}}$ to be reversible, which means $\pi_{i}(\text{d}x_{i}|\textbf{x}_{-i})P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})=P_{i}^{\textbf{x}_{-i}}(y_{i},\text{d}x_{i})\pi(\text{d}y_{i}\mid\textbf{x}_{-i})$, and positive semi-definite, which means that $\int_{\mathcal{X}_{i}}f(x_{i})P_{i}^{\textbf{x}_{-i}}f(x_{i})\pi_{i}(\text{d}x_{i}|\textbf{x}_{-i})\geq 0$ for every square integrable $f:\mathcal{X}_{i}\,\to\,\mathbb{R}$, where $P_{i}^{\textbf{x}_{-i}}f(x_{i})=\int f(y_{i})P_{i}^{\textbf{x}_{-i}}(x_{i},\text{d}y_{i})$. The latter are common assumptions which are often satisfied: for example every reversible operator $Q$ can be made positive semi-definite by considering its lazy version, i.e. $\tilde{Q}=\frac{1}{2}I+\frac{1}{2}Q$ with $I$ being the identity operator [39].
@@
$\kappa\left(P_{i}^{\textbf{x}_{-i}}\right)=\inf\left\{\frac{\int_{B}P^{\textbf{x}_{-i}}_{i}(x_{i},B^{c})\pi_{i}(dx_{i}|\textbf{x}_{-i})}{\pi_{i}(B\mid\textbf{x}_{-i})\pi_{i}(B^{c}\mid\textbf{x}_{-i})};\,\pi_{i}(B\mid\textbf{x}_{-i})\in(0,1),B\subset\mathcal{X}_{i}\right\}$ $\textbf{x}_{-i}\in\mathcal{X}_{-i}\,.$ (8)
The latter measures how much the invariant update of $P_{i}^{\textbf{x}_{-i}}$ is close to exact sampling as in $G_{i}^{\textbf{x}_{-i}}$. In particular, one has $\kappa\left(G_{i}^{\textbf{x}_{-i}}\right)=1$ for all $i$ and $\textbf{x}_{-i}$, while $\kappa\left(P_{i}^{\textbf{x}_{-i}}\right)\leq 1$ for general reversible and positive-definite updates (see e.g. Corollary 1 below). Thus, $\kappa(P,K)$ measures how much the coordinate updates of $P$ are close to exact sampling as in $G$, over the set $K$. By construction, $\kappa(G,K)=1$ for all $K$.
@@
$\kappa_{i}(P_{i},\mathcal{X})$ should not be confused with the conductance of $P_{i}$ on $\mathcal{X}$. Indeed the latter is always equal to 00, since $P_{i}$ leaves $\textbf{x}_{-i}$ unchanged. Similarly, $\kappa(P,\mathcal{X})$ is not the conductance of $P$ on $\mathcal{X}$, since it only measures the quality of the conditional updates, not directly the convergence speed of $P^{t}$ to $\pi$.
Remark 4.
Equations (8) and (4) consider two slightly different versions of conductance, which differ in the denominator of the inner formulas (e.g. $\pi(A)$ instead of $\pi(A)\pi(A^{c})$ in (4)). This affects the resulting number of at most a factor of $2$ and it allows to avoid unnecessary constants in the theorems to follows, e.g. Corollary 1.
@@
Let $P$ be a $\pi$-invariant coordinate-wise kernel as in (5). For every $A,K\subset\mathcal{X}$ and $i=1,\dots,d$ we have
@@
Moreover, if $P_{i}^{\textbf{x}_{-i}}$ is reversible and positive semi-definite for every $\textbf{x}_{-i}\in\mathcal{X}_{-i}$, then $G_{i}(\partial A)\geq P_{i}(\partial A)$.
@@
When $P_{i}^{\textbf{x}_{-i}}$ is reversible and positive semi-definite, Theorem 1 with $K=\mathcal{X}$ implies
@@
Let $P$ be a $\pi$-invariant coordinate-wise kernel as in (5). Then for every $s\in[0,1/2)$ and $K\subseteq\mathcal{X}$ we have
[…]
with $\kappa(P,K)$ as in (7). Moreover, if $P_{i}^{\textbf{x}_{-i}}$ is reversible and positive semidefinite for every $i=1,\dots,d$ and $\textbf{x}_{-i}\in\mathcal{X}_{-i}$, then $\Phi_{s}(G)\geq\Phi_{s}(P)$.
@@ 4 Applications to Metropolis-within-Gibbs schemes
MwG schemes are an instance of coordinate-wise kernels and thus (9) applies to them. We now consider two instances of MwG schemes where the conditional conductance can be controlled, namely independent MH conditional updates and conditionally log-concave target distributions.
@@ 4.1 Conditional updates with independent Metropolis-Hastings
for some $Q_{i}^{\textbf{x}_{-i}}\in\mathcal{P}(\mathcal{X}_{i})$. Despite its simplicity, MwG with IMH proposals is routinely used in various contexts (see e.g. the Bayesian inference for diffusions example of Section 8 [61, 6, 47]). The next proposition shows that an upper bound on the Radon-Nykodym derivative between $\pi_{i}(\cdot\mid\textbf{x}_{-i})$ and $Q_{i}^{\textbf{x}_{-i}}$, i.e.
@@
While it is well-known that IMH suffers from the curse of dimensionality (i.e. performances usually deteriorate exponentially with $d$ if the update is jointly applied to all coordinates), Proposition 1 implies that MwG schemes with IMH proposals only pay a constant slowdown relative to exact Gibbs if the dimensionality of each $\mathcal{X}_{i}$ is fixed. This result will be used in Section 8 to provide a lower bound on the conductance of a data augmentation scheme for discretely observed diffusions.

### v2

@@ Abstract
We study general coordinate-wise MCMC schemes (such as Metropolis-within-Gibbs samplers), which are commonly used to fit Bayesian non-conjugate hierarchical models. We relate their convergence properties to the ones of the corresponding (potentially not implementable) random scan Gibbs sampler through the notion of conditional conductance. This allows us to study the performances of popular Metropolis-within-Gibbs schemes for non-conjugate hierarchical models, in high-dimensional regimes where both number of datapoints and parameters increase. Given random data-generating assumptions, we establish dimension-free convergence results, which are in close accordance with numerical evidences. Applications to Bayesian models for binary regression with unknown hyperparameters and discretely observed diffusions are also discussed. Motivated by such statistical applications, auxiliary results of independent interest on approximate conductances and perturbation of Markov operators are provided.
@@ 1 Introduction
Over 30 years ago, the Gibbs sampler (GS) and more general coordinate-wise samplers (often termed Metropolis-within-Gibbs (MwG) samplers) were introduced as powerful tools to enable Bayesian inference for structured data (featuring spatial, hierarchical or temporal dependence), for example see [25, 73, 7, 14, 12]. The substantial impact of these methods has spread across a wide variety of scientific fields, for example see [26, 30, 48].
Coordinate-wise Markov chain Monte Carlo (MCMC) schemes work by partitioning the vector of parameters in different blocks and updating them one at the time, conditional on all the others. Working in such a coordinate-wise manner can be computationally beneficial in many cases (see e.g. Section 1.1) and it has been observed empirically [56, 46] that such samplers often have extremely good convergence properties. However theoretical understanding of this phenomenon has proved difficult, as the analysis of the resulting algorithms is often subtle and case-specific. Substantial progress has now been made on understanding the pure GS (see Section 1.3), but implementation of these algorithms is generally restricted to contexts with specialised conditional conjugacy properties and therefore it is important to understand more general coordinate-wise samplers which remain comparatively under-studied.
@@ 1.1 Motivating example: non-conjugate hierarchical models
with $y\in\{0,\dots,m\}$ and $m$ being a positive integer. Thus, conditional on $\theta_{j}$, for each group $j$ a sequence of $m$ independent Bernoulli experiments are performed with success probability $e^{\theta_{j}}/(1+e^{\theta_{j}})$. Under model (1), the conditional distribution of $\bm{\theta}$ given $(Y_{1:J},\mu,\tau)$ factorizes as $\mathcal{L}(\text{d}\bm{\theta}|Y_{1:J},\mu,\tau)=\otimes_{j=1}^{J}\mathcal{L}(\text{d}\theta_{j}|Y_{j},\mu,\tau)$, where $\otimes$ denotes the product of independent distributions. This makes model (1) particularly well-suited for coordinate-wise samplers, since the high-dimensional update of $\bm{\theta}$ given $(\mu,\tau)$ decouples into $J$ independent low-dimensional updates (see Section 6.4 for more discussion on the computational implications of this). Figure 1 compares the efficiency of the resulting samplers with two gradient-based MCMC methods. The target is the joint posterior distribution $\mathcal{L}(\text{d}\bm{\theta},\text{d}\mu,\text{d}\tau|Y_{1:J})$, and we consider the high-dimensional regime where $J\to\infty$, so that both the number of datapoints and parameters, i.e. $n=Jm$ and $p=J+2$ respectively, diverge. Full details on the simulation set-up, including algorithmic and prior specifications, are postponed to Section 6.5.
@@ 1.2 Objective and structure of the paper
The structure of the paper is as follows. Section 2 briefly recalls the notion of conductance and its connection to mixing times of Markov chains. Section 3 introduces the class of coordinate-wise schemes that we consider (which include random scan GS and MwG) and provides general results relating the conductance of coordinate-wise schemes to the one of GS (Theorem 1 and Corollary 1). Section 4 consider the specific case of MwG schemes and the example of conditionally log-concave targets. After discussing independent results about approximate conductances and perturbations of Markov operators (Section 5), we move to statistical applications, where we consider Bayesian hierarchical models (Section 6), logistic regression with unknown hyperparameters (Section 7) and inference for discretely observed diffusion processes (Section 8). Mathematical proofs, together with some additional examples and results, are postponed to the Appendix. R code to replicate all the numerical experiments can be found at https://github.com/gzanella/Metropolis-within-Gibbs/.
@@ 1.3 Related literature
Compared to other MCMC schemes (such as gradient-based ones), there are relatively few quantitative theoretical results for coordinate-wise sampling schemes. Moreover, most works focus on the exact Gibbs sampler, e.g. [72, 1, 66, 41, 37, 80, 58, 34, 59, 62, 15], for which specific techniques exploiting the exact updating structure and the associated alternating projection representation can be exploited [65, 18, 19, 4, 3]. Results for general coordinate-wise MCMC, including MwG, are instead quite limited. Exceptions include [71, 49, 35, 36, 60], which mostly focus on geometric or uniform ergodicity, and [77], which provides some analysis of MALA-within-Gibbs schemes.
While working on our manuscript, it came to our attention that [61] independently and concurrently developed results relating the spectral Gap of coordinate-wise samplers to the ones of exact GS. Their results are similar in spirit to the ones we develop in Sections 3 and 4. On the other hand, the specific results and the type of applications are significantly different. For example, [61] study the spectral gap, while in this work we work with the approximate version of the conductance defined in (4), which is crucial for the subsequent applications we consider (see e.g. Remarks 9 and 12). Also, we consider and develop in some detail various statistical applications (Sections 6, 7 and 8) where we combine our techniques with posterior asymptotics, dimensionality reduction and perturbation arguments.
@@ 2 Mixing times and conductance
In order to give explicit quantitative bounds on (3), we will assume that $P$ is $\pi$-reversible, i.e. $\pi(\text{d}\textbf{x})P(\textbf{x},\text{d}\textbf{y})=\pi(\text{d}\textbf{y})P(\textbf{y},\text{d}\textbf{x})$, and positive semi-definite, meaning that
$\int\left(Pf(\textbf{x})\right)f(\textbf{x})\pi(\text{d}\textbf{x})\geq 0,\quad Pf(\textbf{x})=\int f(\textbf{y})P(\textbf{x},\text{d}\textbf{y}),$
for every $f\,:\,\mathcal{X}\,\to\,\mathcal{X}$ such that $\int f^{2}(\textbf{x})\pi(\text{d}\textbf{x})<\infty$. The latter are common assumptions, which are often satisfied in practice. A typical strategy to determine the convergence properties of a reversible and positive semi-definite Markov chain is to study the $s$-conductance of $P$, i.e.
@@
Let $P$ be a $\pi$-reversible and positive semi-definite Markov transition kernel. For every $s\in[0,1/2)$, $t\geq 0$, $M\geq 1$ and $\mu\in\mathcal{N}\left(\pi,M\right)$, it holds
@@
The usual definition of $s$-conductance (see e.g. [45, 21]) is slightly different with respect to the one given in (4). Appendix C provides results for the alternative definition. Definition (4) is less tight but leads to neater formulas (see e.g. Theorem 1).
@@ 3 Coordinate-wise MCMC
where $P_{i}^{\textbf{x}_{-i}}$ is a $\pi_{i}(\cdot|\textbf{x}_{-i})$-invariant Markov kernel on $\mathcal{X}_{i}$ and thus, by construction, $P_{i}$ is invariant with respect to $\pi$. At each iteration, a Markov chain evolving according to $P$ picks uniformly at random a coordinate $i$ in $\{1,\dots,d\}$ and then updates $x_{i}$ sampling from $P_{i}^{\textbf{x}_{-i}}(x_{i},\cdot)$, while leaving $\textbf{x}_{-i}$ unchanged. Lemma 5 in the Appendix shows that reversibility and positive semi-definiteness of $P$ are implied by the ones of $P_{i}$, which are in turn equivalent to $\pi(\cdot\mid\textbf{x}_{-i})$-reversibility and positive semi-definiteness of $P_{i}^{\textbf{x}_{-i}}$ for almost every $\textbf{x}_{-i}$. Thus, in the following we will always silently assume that $P_{i}^{\textbf{x}_{-i}}$ is reversible and positive semi-definite, for every $i$ and $\textbf{x}_{-i}$. The latter are common assumptions which are often satisfied: for example they hold for the operators $P^{\textbf{x}_{-i}}$ defined in Section 4.2. Moreover, every reversible operator $Q$ can be made positive semi-definite by considering its lazy version, i.e. $\tilde{Q}=\frac{1}{2}I+\frac{1}{2}Q$ with $I$ being the identity operator [45].
@@
Remark 3.
Another popular version of coordinate-wise methods, called deterministic scan, updates the coordinates according to a pre-specified order. The relationship between different deterministic and random orders is not univocal: even if we expect them to behave in a similar way in most cases, for some specific problems the associated mixing times can significantly vary (see [67, 31, 23]). In this work we restrict to random scan schemes, which somewhat simplify the mathematical treatment: for example, random scan Gibbs as in (6) is always positive semi-definite.
@@
$\kappa\left(P_{i}^{\textbf{x}_{-i}}\right)=\inf\left\{\frac{\int_{B}P^{\textbf{x}_{-i}}_{i}(x_{i},B^{c})\pi_{i}(dx_{i}|\textbf{x}_{-i})}{\pi_{i}(B\mid\textbf{x}_{-i})};\,\pi_{i}(B\mid\textbf{x}_{-i})\in\left(0,\frac{1}{2}\right],B\subset\mathcal{X}_{i}\right\}$ $\textbf{x}_{-i}\in\mathcal{X}_{-i}\,.$ (8)
The latter measures how much the invariant update of $P_{i}^{\textbf{x}_{-i}}$ is close to exact sampling as in $G_{i}^{\textbf{x}_{-i}}$.
@@
$\kappa_{i}(P_{i},\mathcal{X})$ should not be confused with the conductance of $P_{i}$ on $\mathcal{X}$. Indeed the latter is always equal to 0, since $P_{i}$ leaves $\textbf{x}_{-i}$ unchanged. Similarly, $\kappa(P,\mathcal{X})$ is not the conductance of $P$ on $\mathcal{X}$, since it only measures the quality of the conditional updates, not directly the convergence speed of $P^{t}$ to $\pi$.
@@
Let $P$ be as in (5). For every $A,K\subset\mathcal{X}$ and $i=1,\dots,d$ we have
@@
Moreover, $G_{i}(\partial A)\geq P_{i}(\partial A)$.
@@
Theorem 1 with $K=\mathcal{X}$ implies
@@
Let $P$ be as in (5). Then for every $s\in[0,1/2)$ and $K\subseteq\mathcal{X}$ we have
$\Phi_{s}(P)$ $\geq\kappa(P,\mathcal{X})\Phi_{s}(G)\quad\text{and}\quad\Phi_{s}(P)\geq\kappa(P,K)\Phi_{s}(G)-\frac{\pi(K^{c})}{s}\left(\frac{1}{d}\sum_{i=1}^{d}\kappa_{i}(P_{i},K)\right)\,$, (9)
with $\kappa(P,K)$ as in (7). Moreover, $\Phi_{s}(G)\geq\Phi_{s}(P)$.
@@
Remark 7.
The results of Corollary 1 extend to the case where the update probabilities in (5) are not uniform. See Corollary 3 in the Appendix for a precise statement.
@@ 4 Applications to Metropolis-within-Gibbs schemes
MwG schemes are an instance of coordinate-wise kernels and thus (52) applies to them. Moreover, it is well-known that any operator as in (10) is reversible (e.g. Theorem $7.2$ in [63]). We now consider two instances of MwG schemes where the conditional conductance can be controlled, namely independent MH conditional updates and conditionally log-concave target distributions.
@@ 4.1 Conditional updates with independent Metropolis-Hastings
for some $Q_{i}^{\textbf{x}_{-i}}\in\mathcal{P}(\mathcal{X}_{i})$. Despite its simplicity, MwG with IMH proposals is routinely used in various contexts, e.g. Bayesian inference for diffusions example of Section 8 [69, 9, 54]. Moreover, a MwG operator as in (11) is always positive semi-definite, see Lemma $3.1$ in [5]. The next proposition shows that an upper bound on the Radon-Nykodym derivative between $\pi_{i}(\cdot\mid\textbf{x}_{-i})$ and $Q_{i}^{\textbf{x}_{-i}}$, i.e.
@@
While it is well-known that IMH suffers from the curse of dimensionality (i.e. performances often deteriorate exponentially with $d$ if the update is jointly applied to all coordinates, see e.g. Figure 3, or [17, Prop.2] for lower bounds on IMH asymptotic variances), Proposition 1 implies that MwG schemes with IMH proposals only pay a constant slowdown relative to exact Gibbs if the dimensionality of each $\mathcal{X}_{i}$ is fixed. This result will be used in Section 8 to provide a lower bound on the conductance of a data augmentation scheme for discretely observed diffusions.